OpenStudy (anonymous):

Solve the word problem. The table shows the low outside temperatures for Monday, Tuesday, and Wednesday. By how many degrees did the low temperature increase from Monday to Tuesday? Addition and Subtraction Applications A. 5.6° B. 11.9° C. 17.5° D. 23.1°

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

the requested change is: \(-5.6-(-17.5)= -5.6+17.5=... \)